# Training Noise Token Pruning

### v1

@@ 1 Introduction
Token pruning methods exploit predictive information differences between tokens; for a given task, some tokens are more useful than others. By dropping the least informative tokens first, a learned method can preserve the most amount of accuracy while removing desired number of tokens. While these methods were originally explored in a natural language processing context, token redundancy is arguably stronger in image transformers, and multiple token pruning methods for vision models (e.g. ViT [8]) have been proposed [17, 23, 21, 22, 15]. Unfortunately, token pruning as an action is discrete and not easily optimized directly. Current techniques use either forms of stochastic discrete dropping [17, 23], or heuristics [21] using attention activations themselves.
In this paper, we reframe Token Pruning as a constrained case of the Information Bottleneck problem. By viewing the token dropping rate as a channel constraint and the accuracy penalty as a distortion metric, we can directly apply ideas from the compression literature. Relaxations of these conditions provide continuous optimization desiderata, which are more easily optimized. Even though the original problem is difficult to solve exactly due to high-order token interaction information (“synergy” and “redundancy”), we show that a simple approximation to this case works surprisingly well. Specifically, we demonstrate that the proposed method has superior performance in comparison to recent pruning methods using either stochastic discrete dropping or attention heuristics on standard base models for the ImageNet-1K benchmark, in particular in the low token regime (i.e., when removing a large percentage of tokens).
@@ 3 Methodology
3.1 Token pruning as bottleneck optimization
We model token pruning optimization as a set selection problem. Given a transformer architecture $f$, we want to select the token subset $T\subset\mathcal{T}$ that minimizes loss $\mathcal{L}$ for each datapoint and label pair $(x,y)$ subject to constraint $|T|<K$:
$\min_{T\subset\mathcal{T}}\mathcal{L}(f(x|_{T}),y)\text{ s.t. }|T|<K$ (1)
where $\mathcal{T}$ is the set of all tokens, and $x|_{T}$ are input data restricted to the token set $T$.
We can map this problem into the information-theoretic literature by drawing analogies between the loss and constraint functions in Eq. 1 and the rate and distortion functions of compression literature [18]. The loss function usually measures discrepancies between $y$ and its estimator $f(x|_{T})$, and, as in previous work [18, 2], is a direct analog to the distortion function.
The analog to the rate function requires more assumptions. While viewing each token as a fixed number of bits directly maps the $|T|<K$ constraint into a transmission rate, this does nothing to aid our optimization. Taking the mapping one step further, we could instead use the mutual information $I(x|_{T},x)<K$ constraint. This forces the effective information transmission to remain below $K$ bits. We then would like to solve
$\min_{T\subset\mathcal{T}}\text{Distortion}(f(x|_{T}),y)\text{ s.t. }I(x|_{T},x)<K$ (2)
which is exactly the Information Bottleneck.
This formulation is at once quite useful in that it maps a discrete optimization into a continuous problem, yet unfortunately also problematic in its exact computation. The rate parameter is definitionally aware of the content of the tokens. This provides us with a route to a potential procedure: force the token content to have at most $K$ bits of information from the original $x$ by adding an external source of noise. Selecting what information to keep from $x$ is equivalent to token selection, but still enjoys a continuous optimization landscape, so long as we choose to add continuous noise. On the otherhand, in order to exactly measure the mutual information we need to measure redundancy between tokens. This issue also was implicitly present in the discrete form of the optimization problem, but it is explicitly included in mutual information.
3.2 Proposed approximate solution
Solutions to the information bottleneck have been proposed in the literature, both in a general setting [18] and in a deep learning context [1, 2]. However, for token pruning, these solutions are almost certainly too computationally heavy to be efficient as dropping methods. Instead, we will solve a simplified version that ignores redundant information, then prune redundant tokens in a second phase. Surprisingly, even with the errors induced by approximation, this method produces accurate results with very little compute overhead.
Given a non-pruned vision transformer architecture, we train a module we call the Noise Allocator to estimate relative token relevance scalars $\alpha_{i}$ for each token. Due to the residual connections used by many transformer architectures, this module can be placed after any of the individual transformer blocks, or after multiple blocks (in a multi-layer dropping case).
Consider an architecture with L repeated transformer blocks possibly after an initial token embedding layer (for example, ViT has 12 identical blocks). Let the $\ell$th Transformer block $f^{(\ell)}$ with input $x\in\mathbb{R}^{N\times D}$, where $N$ is the number of tokens, and $D$ is the embedding dimension. Let $x^{(\ell)}_{i}$ denote the $i$th token. For every block where the user would like to perform pruning, we introduce a linear token relevance predictor with a weight matrix $W^{(\ell)}\in\mathbb{R}^{D\times 1}$.
We then compute $\alpha_{i}$ as
@@ 3.2 Proposed approximate solution
For training/fine-tuning, we compute noise variables $\eta_{i}^{\ell}$ as
@@
where $\varepsilon\sim\mathcal{N}(0,\beta I_{D})$ (“the reparameterization trick” [14]). Here, $\beta$ is a hyper-parameter that controls the total amplitude of noise added, and corresponds with $K$ up to a monotonic transformation. The Softmax function in Eq. 3 ensures that the network must allocate $\beta$ amount of noise; without it, a degenerate zero-noise solution exists.
Classical results in Information Theory state that the mutual information (channel capacity) for each token to its noised variant has an upper bound proportional to $\log(1+P_{signal}/P_{noise})$, where $P_{signal}/P_{noise}$ is the ratio of the power (amplitude) of the token signal and to that of the noise [6] (the signal to noise ratio). Due to the layer-wise normalization (LayerNorm) of many transformer architectures, $P_{signal}$ is necessarily bounded. Disregarding the synergistic information between tokens, optimizing our proposed Noise Allocator amounts to solving an information bottleneck problem. We discuss limitations of this approach in Section 4.4.
At test time noise is not sampled, and instead of $\beta$ we directly select the top $K$ tokens, dropping the rest of the token set. We interpret $\alpha^{(\ell)}$ as a token importance map for $x^{(\ell)}$. Based on this importance ranking, we may either discard less important tokens or merge them with other tokens in the subsequent stage.
Weight vector $W$ is applied to every $x_{i}$ identically, so the additional parameter space complexity remains $\mathcal{O}(NL)$, where $L$ is the number of layers at which tokens are selectively dropped during testing. Given $L\ll D$, this complexity effectively reduces to $\mathcal{O}(D)$, making it comparatively negligible within the overall model size. For weaker encoders and/or smaller transformer blocks such as DeiT-Tiny, it is sometimes helpful to increase the complexity of $W$ to a multi-layer structure (e.g., an MLP), as we show in experiments in the Appendix.
3.2.1 Similarity-based Pruning by Random Partition (Redundancy Removal)
Even though individual tokens might be highly relevant to the classification problem (i.e, individually $I(x_{i}^{\ell},y)$ might be high), relevant tokens sharing a high amount of information with a kept token represents wasted capacity. In order to avoid this waste, we implement a redundancy removal step for test time usage, which removes similar and therefore possibly redundant tokens.
@@ 3.2.1 Similarity-based Pruning by Random Partition (Redundancy Removal)
Figure 4: Single Layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for single layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column.
K=1.0 K=0.8 K=0.6 K=0.5 K=0.25
#tokens=196 #tokens=156 #tokens=127 #tokens=98 #tokens=49
Random Drop 80.50 4.63 79.54 3.90 77.83 3.20 76.57 2.87 67.77 2.03
ToMe [5] 80.07 3.85 77.40 3.11 73.13 2.75 - -
Zero-TP [21] 80.00 3.91 78.94 3.21 77.74 2.88 69.99 2.05
DynamicViT [17] 79.39 3.99 78.66 3.28 77.96 2.94 71.03 2.08
Top-K [22] 80.12 3.85 78.89 3.11 77.69 2.75 68.57 1.86
TNT (ours) 80.11 3.90 79.29 3.20 78.65 2.87 72.66 2.03
#tokens=196 #tokens=160 #tokens=130 #tokens=100 #tokens=60
ViT/16 Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs
Random Drop 78.70 9.17 77.17 7.69 75.12 6.50 69.21 5.33 31.16 3.81
ToMe [5] 77.71 7.66 75.70 6.42 66.46 5.22 - -
Zero-TP [21] 77.43 7.32 75.96 6.52 70.92 5.35 40.84 3.83
TNT (ours) 77.94 7.70 76.22 6.50 71.69 5.33 41.75 3.81
Figure 5: Multi-layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for multi-layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column.
@@ 4 Experiments
We conduct a series of experiments on pre-trained vision transformer models, including DeiT (Tiny [^1] DeiT-Tiny results relegated to the Appendix., Small, Base) [19] and ViT/16 [8], each trained on ImageNet-1K [7]. We then compare our model with previously proposed token-dropping models, as well as two simple baseline schemes: DynamicViT [17], ZeroTP[21], ToMe [5], Top-K [11], and random dropping. In Section 4.3, we present ablation studies to validate our individual design choices.
@@ 4.1 Rate Sweep for Single Layer Pruning
The experiments of previous studies have shown that most models [17, 9, 21] refrain from dropping tokens in initial Transformer layers, as these layers contain limited information relevant to token importance. For our experiments, we select the second layer for ViT and third layer for DeiT variants as the standard layers for single-layer token pruning. We use the standard pre-trained DeiT variants for all experiments for all models. For all ViT experiments for all models, we use a modified ViT/16 with smaller embedding dimension $mlp\_dim=emb\_dim=768$. Additionally, we use mean-pooled token embeddings for predictions instead of the CLS token.
@@ 4.2 Rate Sweep at Multi-layer Pruning
We evaluate the performance of token pruning across multiple layers as this better reflects a performant pruning system, with the caveat that equal comparison conditions are more difficult to ensure. For the token keep rate at each layer and the specific layers chosen for token pruning (i.e. pruning locations), we strictly follow the instructions on the original paper for each model. For Zero-TP [21], tokens are pruned at layers [1, 3, 6, 9, 11], where layers [1, 11] perform similarity-based pruning only; for ToMe [5], pruning occurs at every layer; and for DynamicViT [17], pruning is applied at layers [3, 6, 9]. As Top-K lacks specific pruning instructions, we align its pruning layers with those of TNT. We gradually sweep the keep rate at pruning locations to generate Figure 5 for DeiT-B-Distil., DeiT-S-Distil., and ViT/16. We also provide partial numerical results in Table 2. All base models are the same as those used in the single-layer pruning section. Additional details on experimental setups and specific details for each model are provided in the Appendix.
For TNT, we prefer pruning tokens at earlier layers, as it provides stronger performance, and as shown in Figure 3, layers [3, 4, 5] serve as effective locations for token pruning. Rather than performing the similarity pruning stage at every pruning layer as in Zero-TP, we prune $s$ tokens before tokens are sent to the ViT blocks. Our ablation study justify this. For all experiments, $s$ is set to 40. Our results indicate that TNT consistently shows strong performance in varying token (computation) ranges. As the proportion of pruned tokens increases, the importance of the retained tokens becomes more critical. TNT shows a larger performance gap compared to other models in the low-token regime.
@@ 4.4 Limitations
While the proposed TNT method provides relatively improved performance over other methods, important gaps remain in methodology. In this method, redundant tokens are not removed during training; while this can be performed through merging as in ToMe [5], as experimental results show it is difficult to do in a stable manner. Further, synergistic information between tokens is not considered; this seems less relevant to ImageNet classification, but likely would be useful in more complicated label structures (e.g., hierarchical/multi-class structures). This is also perhaps a deeper problem than redundant tokens, as it has combinatorial complexity in the order of the interactions considered.

### v2

@@ 1 Introduction
Token pruning methods exploit predictive information differences between tokens; for a given task, some tokens are more useful than others. By dropping the least informative tokens first, a learned method can preserve the most amount of accuracy while removing desired number of tokens. While these methods were originally explored in a natural language processing context, token redundancy is arguably stronger in image transformers, and multiple token pruning methods for vision models (e.g. ViT [7]) have been proposed [16, 22, 20, 21, 14].
Token relevance in many transformer models is already captured by the CLS token, and the attention values it attributes to the other tokens. The CLS tokens are directly predictive of their given task, often trained using the appropriate prediction loss, mapped to the label domain by a single linear layer. Thus, the tokens that the CLS token attends to are the relevant tokens for that task. This intuition is implemented in [14, 8]; moreover, Haurum et al. [10] show that simply ranking tokens by CLS attention score and taking the desired number (the “Top-K” method) outperforms most methods proposed to date. However, these methods are conditional on the CLS token existing in the architecture, and the primary task matching the CLS token training task.
In this paper, we provide high fidelity estimates of relevance without the CLS token by using a variational Information Bottleneck [17, 1]. The VIB optimization objective is a trade-off between transmission rate (the number of bits in a message) and relevance to some extrinsic factor. This is exactly the same trade-off in Token Dropping, where we would like minimal token counts (analogous to transmission rate) for maximal relevance. We demonstrate that our proposed method has superior performance in comparison to recent pruning methods using either stochastic discrete dropping and matches or exceeds the performance of CLS-attention methods even without the CLS token, both in terms of accuracy and in computational cost.
@@ 3 Methodology
The Information Bottleneck [17] is a Rate-Distortion trade-off problem for transmitting a signal $s$ relevant to labels $y$ through an encoding $x$. In its setup, we’re asked to find the minimal transmission rate, i.e., the minimal effective number of bits sent to $x$, which is the mutual information $I(s,x)$, subject to a maximal distortion of $x$’s relevance to $y$, which is measured by $I(x,y)$. This is written as a constrained optimization
$\min_{p(z|x)}I(s,x)\text{ subject to }I(x,y)>T.$ (1)
This is often rewritten as an unconstrained optimization
$\min_{p(z|x)}I(s,x)-\lambda I(x,y)$ (2)
where $\lambda$ is a function of the original constraint $T$. Dividing by $\lambda$, there is an equivalent objective $\frac{1}{\lambda}I(s,x)-I(x,y)$, we can then rewrite the optimization as a maximization of relevance (or a minimization of negative relevance) constrained by effective transmission rate:
$\min_{p(z|x)}-I(x,y)\text{ subject to }I(s,x)<\tilde{T}.$ (3)
While Eq. 3 is not the original formulation of the bottleneck objective and is arrived at using elementary algebraic manipulations, it more clearly illustrates the analogy to the Token Pruning problem. We wish to limit the number of tokens transmitted (and subsequently processed, etc.) to some fixed number L, while maintaining the highest prediction accuracy, which is equivalent to minimizing the cross entropy $H(y|x)$, which itself is the objective $-I(x,y)$ up to constant $-H(y)$.
We can approximate solutions to the information bottleneck using the deep variational method provided by Alemi et al. [1]. Consider an architecture with L repeated transformer blocks possibly after an initial token embedding layer (for example, ViT has 12 identical blocks), where the $\ell$th Transformer block $f^{(\ell)}$ with input $x\in\mathbb{R}^{N\times D}$ has $N$ tokens with embedding dimension $D$, and individual $x^{(\ell)}_{i}$ tokens. For every block where we would like to perform pruning, we introduce a linear token relevance predictor with a weight matrix $W^{(\ell)}\in\mathbb{R}^{D\times 1}$.
We then compute $\alpha^{(\ell)}_{i}$ as
@@
We then compute noise variables $\eta_{i}^{\ell}$ as
@@
where $\varepsilon\sim\mathcal{N}(0,\beta I_{D})$ (“the reparameterization trick” [13]). Here, $\beta$ is a hyper-parameter that controls the total amplitude of noise added. The $\alpha^{(\ell)}_{i}$’s are our estimate of the most relevant tokens; if the network could output arbitrary values in [0,1] for $\alpha$, clearly the highest accuracy solution sends all $\alpha^{(\ell)}_{i}$’s to 1, adding zero noise and thus preserving all the signal. Due to the Softmax, instead this noise allocation is constrained to always add $\beta$ amount of noise. This forces some tokens to be dropped; the solution that has the highest prediction accuracy is the one that adds the least noise (i.e., has the highest $\alpha^{(\ell)}_{i}$) to the most predictive tokens. Thus, $\alpha^{(\ell)}_{i}$ is an approximation to the relevance of $x^{(\ell)}_{i}$.
At test time, instead of adding noise, we use the $\alpha^{(\ell)}_{i}$ to rank tokens in order of relevance. By removing all but the top ranked solutions (similar to what the “Top-K” method does for the CLS token attention scores).
Classical results in Information Theory state that the mutual information (channel capacity) for each token to its noised variant has an upper bound proportional to $\log(1+P_{signal}/P_{noise})$, where $P_{signal}/P_{noise}$ is the ratio of the power (amplitude) of the token signal and to that of the noise [5] (the signal to noise ratio). Due to the layer-wise normalization (LayerNorm) of many transformer architectures, $P_{signal}$ is necessarily bounded. This solution can also be directly mapped onto the Deep Variational Information Bottleneck by viewing our initial tokens $x^{(\ell)}_{i}$ as the mean of their latent embeddings, and our $(1-\alpha^{(\ell)}_{i})$ as the element-wise standard deviations.
3.1 Similarity-based Pruning by Random Partition (Redundancy Removal)
Our chosen bottleneck approximation only considers elementwise approximation. Both $I(x,y)$ and $I(s,x)$ measure interaction information; conditional on adding one token $x_{i}$, the information about $y$ in another token $x_{j}$ might be higher than the marginal information $I(x_{j},y)$. These synergies are not limited to dyads, and can have arbitrarily high order. Further, redundancies, where $I(x_{i},y)+I(x_{j},y)>I((x_{i},x_{j}),y)$ mean that we might include additional, unnecessary information. Even though individual tokens might be highly relevant to the classification problem (i.e, individually $I(x_{i}^{\ell},y)$ might be high), relevant tokens sharing a high amount of information with a kept token represents wasted capacity. This weakness is shared by CLS-token based removal methods, and in order to avoid this waste, we implement the same redundancy removal method. We discuss its limitations in Section 4.4.
@@ 3.1 Similarity-based Pruning by Random Partition (Redundancy Removal)
Figure 4: Single Layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for single layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column. Note that the mean-pooled token embedding ViT in the third column has no CLS token, and thus EViT and Top-K cannot be applied to it.
@@
K=1.0 K=0.8 K=0.6 K=0.5 K=0.25
#tokens=196 #tokens=156 #tokens=127 #tokens=98 #tokens=49
Random Drop 80.50 4.63 79.54 3.90 77.83 3.20 76.57 2.87 67.77 2.03
ToMe [4] 80.07 3.85 77.40 3.11 73.13 2.75 - -
Zero-TP [20] 80.00 3.91 78.94 3.21 77.74 2.88 69.99 2.05
DynamicViT [16] 79.39 3.99 78.66 3.28 77.96 2.94 71.03 2.08
Top-K [10] 80.12 3.85 78.89 3.11 77.69 2.75 68.57 1.86
EViT [14] 78.77 3.86 76.07 3.27 76.28 2.74 69.67 1.86
TNT (ours) 80.11 3.90 79.29 3.20 78.65 2.87 72.66 2.03
#tokens=196 #tokens=160 #tokens=130 #tokens=100 #tokens=60
ViT/16 Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs Acc. GFLOPs
Random Drop 78.70 9.17 77.17 7.69 75.12 6.50 69.21 5.33 31.16 3.81
ToMe [4] 77.71 7.66 75.70 6.42 66.46 5.22 - -
Zero-TP [20] 77.43 7.32 75.96 6.52 70.92 5.35 40.84 3.83
DynamicViT [16] 77.18 7.98 75.25 6.68 70.77 5.43 39.68 4.06
TNT (ours) 77.94 7.70 76.22 6.50 71.69 5.33 41.75 3.81
Figure 5: Multi-layer Pruning results: We plot the Top-1 Accuracy in the ImageNet-1k validation set for each of the pruning methods as a function of computational efficiency, in the top row measured by GFLOPs and in the bottom row measured by throughput, for multi-layer pruning. The base model is DeiT-B-Distil in the first column, DeiT-S-Distil. in the second column, and ViT/16 in the third column. Note that the mean-pooled token embedding ViT in the third column has no CLS token, and thus EViT and Top-K cannot be applied to it.
@@ 4 Experiments
We conduct a series of experiments on pre-trained vision transformer models, including DeiT (Tiny [^1] DeiT-Tiny results relegated to the Appendix., Small, Base) [18] and ViT/16 [7], each trained on ImageNet-1K [6]. We then compare our model with previously proposed token-dropping models, as well as two simple baseline schemes: DynamicViT [16], ZeroTP[20], ToMe [4], Top-K [10], EViT[10] and random dropping. Because ATS [8] adaptively prunes a variable number of tokens per image, it does not allow a systematic sweep of the keep rate. Therefore, we do not include it as a baseline. We also validate our individual design choices in an ablation study (Section 4.3).
Haurum et al. [10] identifies CLS token attention based methods as state-of-the-art [10, 8], with a naive baseline (“Top-K”) nearly outperforming all others. These methods cannot be applied to models without CLS tokens, or where the CLS token is mismatched with the test-time task. However, our proposed method, which operates on similar principles but without the CLS token, can be applied in such settings, as demonstrated by experiments on the mean-pooled token embedding ViT.
@@ 4.1 Rate Sweep for Single Layer Pruning
The experiments of previous studies have shown that most models [16, 8, 20] refrain from dropping tokens in initial Transformer layers, as these layers contain limited information relevant to token importance. For our experiments, we select the second layer for ViT and third layer for DeiT variants as the standard layers for single-layer token pruning. We use the standard pre-trained DeiT variants for all experiments for all models. For all ViT experiments for all models, we use a modified ViT/16 with smaller embedding dimension $mlp\_dim=emb\_dim=768$.
@@ 4.2 Rate Sweep at Multi-layer Pruning
We evaluate the performance of token pruning across multiple layers as this better reflects a performant pruning system, with the caveat that equal comparison conditions are more difficult to ensure. For the token keep rate at each layer and the specific layers chosen for token pruning (i.e. pruning locations), we strictly follow the instructions on the original paper for each model. For Zero-TP [20], tokens are pruned at layers [1, 3, 6, 9, 11], where layers [1, 11] perform similarity-based pruning only; for ToMe [4], pruning occurs at every layer; and for EViT [14] and DynamicViT [16], pruning is applied at layers [3, 6, 9] and [4, 7, 10], respectively. As Top-K [10] lacks specific pruning instructions, we align its pruning layers with those of TNT. We gradually sweep the keep rate at pruning locations to generate Figure 5 for DeiT-B-Distil., DeiT-S-Distil., and ViT/16. We also provide partial numerical results in Table 1. All base models are the same as those used in the single-layer pruning section. Additional details on experimental setups and specific details for each model are provided in the Appendix.
For TNT, we prefer pruning tokens at earlier layers, as it provides stronger performance, and as shown in Figure 3, layers [3, 4, 5] serve as effective locations for token pruning. Rather than performing the similarity pruning stage at every pruning layer as in Zero-TP, we find it is better to prune $s$ tokens before tokens are sent to the ViT blocks. For all experiments, $s$ is set to 40. Our results indicate that TNT consistently shows strong performance in varying token (computation) ranges. As the proportion of pruned tokens increases, the importance of the retained tokens becomes more critical. TNT shows a larger performance gap compared to other models in the low-token regime.
@@ 4.4 Limitations
While the proposed TNT method provides relatively improved performance over other methods, important gaps remain in methodology. In this method, redundant tokens are not removed during training; while this can be performed through merging as in ToMe [4], as experimental results show it is difficult to do in a stable manner. Further, synergistic information between tokens is not considered; this seems less relevant to ImageNet classification, but likely would be useful in more complicated label structures (e.g., hierarchical/multi-class structures). This is also perhaps a deeper problem than redundant tokens, as it has combinatorial complexity in the order of the interactions considered. While the Information Bottleneck technically addresses this, estimating models which include interaction information are generally intractable unless the variables are transformed (which makes them useless for Token Pruning).